# Depth-Optimal Addressing of 2D Qubit Array with 1D Controls Based on Exact Binary Matrix Factorization

### 24 Jan 2024 v1

@@ Abstract
Reducing control complexity is essential for achieving large-scale quantum computing, particularly on platforms operating in cryogenic environments. Wiring each qubit to a room-temperature control poses a challenge, as this approach would surpass the thermal budget in the foreseeable future. An essential tradeoff becomes evident: reducing control knobs compromises the ability to independently address each qubit. Recent progress in neutral atom-based platforms suggests that rectangular addressing may strike a balance between control granularity and flexibility for 2D qubit arrays. This scheme allows addressing qubits on the intersections of a set of rows and columns each time. While quadratically reducing controls, it may necessitate more depth. We formulate the depth-optimal rectangular addressing problem as exact binary matrix factorization, an NP-hard problem also appearing in communication complexity and combinatorial optimization. We introduce a satisfiability modulo theories-based solver for this problem, and a heuristic, row packing, performing close to the optimal solver on various benchmarks. Furthermore, we discuss rectangular addressing in the context of fault-tolerant quantum computing, leveraging a natural two-level structure.
@@ I Introduction
Quantum computing demands a thermal environment with significantly lower fluctuations than the energy level differences in qubits. Because of this, solid-state platforms, such as superconducting circuits or semiconductor quantum dots, can only operate in cryogenic environments. Controlling a quantum processor entirely from room temperature becomes impractical for large-scale quantum computing, where the requirement of millions of physical qubits exceeds the capacity of dilution refrigerators, supporting only hundreds of coaxial cables [1]. Given a certain degree of homogeneity in qubit fabrication, one viable strategy is manipulating multiple qubits with one signal. Hornibrook et al. [2] demonstrate a microarchitecture utilizing ‘switch matrices’ to selectively distribute analog signals. In this work, we extend it to 2D and aim to reduce the control complexity by exploring the idea of relaxing qubit addressability.
Figure 1: Quantum architecture employing the rectangular addressing for a 2D qubit array. a) Conceptual image of the architecture. Analog lines provide the signals to manipulate the qubits. They are first routed by the row switches using the row address signal, and later routed by the qubit switches using the column address signal. The row and column address signals transmit 6 bits per clock cycle, in contrast to 9 bits that would be used for individual addressing. In general, this is a quadratic difference. b) Addressed qubits at different stages of an example quantum circuit. At each stage, only a rectangle of qubits are addressed, which may necessitate more depth, e.g., $t_{4}$ and $t_{5}$ cannot be merged.
For a 2D array $X\times Y$, a (combinatorial) rectangle is a set of the form $X^{\prime}\times Y^{\prime}$, where $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$. Specifying a rectangle requires $|X|+|Y|$ bits (one bit for each row and each column), a significant reduction compared to $|X|\cdot|Y|$ bits for all elements. In Figure 1a, we present a conceptual architecture where qubits can be addressed using rectangles formed by intersections of rows and columns. Compared to a single address line, there are both row and column address lines, and an additional switch for each row. This setting provides an advantage: in each clock cycle, the row and column address lines transmit only the square root of the total bits required in a single address line configuration, at the expense of possible layer execution time. This quadratic reduction is maintained while preserving individual addressability, as a single element can still be considered a rectangle. For example, in Figure 1b, we illustrate qubits being addressed at different stages of a quantum circuit. At $t_{2}$, the column and row addresses are $(0,1,0)$ and $(0,0,1)$ respectively, ensuring only $q_{5}$ is addressed.
Our motivation stems from recent successful large-scale experiments on the neutral atom arrays platform [3] that highlights the effect of reducing control complexity by rectangular addressing. As depicted in Figure 2a, the acousto-optic deflector (AOD) illuminates a product of rows and columns. Quantum gates, induced by specific pulses modulated by the AOD, address qubits at the row and column intersections. AODs prove effective for implementing gates [4, 3] and qubit movements [5, 6, 7, 8]. Moreover, certain key applications, such as the surface code [9], inherently feature rectangular structures. This code requires a 2D grid of data qubits, and another 2D grid of ancilla qubits to conduct error checks. Exploiting this structure, Versluis et al. [10] show that 8 sets of signals are sufficient for the decoding sequence on superconducting qubits. The qubits addressed by each set forms a rectangle, so, in principle, the sequence can be realized by addressing 8 control signals to all qubits within their respective rectangles.
@@
The coarser granularity of rectangular addressing may reduce control complexity at the cost of increasing depth. Although $t_{4}$ and $t_{5}$ in Figure 1b involve single-qubit gates on different qubits, they cannot be combined into a single stage, because merging them would result in a non-rectangular arrangement of qubits to address. A more general problem is given by the matrix in Figure 2b, where the qubits to address are represented by the 1’s. This matrix can be partitioned into five rectangles, each designated by distinct markers. Consecutively, each rectangle receives a modulated $R_{z}$ pulse through specific AOD configurations. Minimizing the number of rectangles to partition arbitrary binary matrices becomes crucial.
In this paper, we consider the problem of achieving depth-optimal rectangular addressing, which we formulate as exact binary matrix factorization. We present a satisfiability modulo theories (SMT, extension of SAT) formulation for this problem and an effective heuristic dubbed row packing. The combined algorithm, SAP (SMT and packing), finds high-quality heuristic solutions quickly and then iteratively approaches to the optimal solution. To assess our methods, we generate three benchmark sets. The first set comprises random matrices, the second includes matrices with known optimal solutions, despite the problem generally being NP-hard. The third set is constructed to accentuate the gap between the partition number and the conventional matrix rank. We observe that for future fault-tolerant quantum computing (FTQC), the problem may exhibit a product structure. This implies that we can solve limited-size problems on multiple levels and then combine the solutions.

### 22 Mar 2024 v2

@@ Abstract
Reducing control complexity is essential for achieving large-scale quantum computing. However, reducing control knobs may compromise the ability to independently address each qubit. Recent progress in neutral atom-based platforms suggests that rectangular (row-column) addressing may strike a balance between control granularity and flexibility for 2D qubit arrays. This scheme allows addressing qubits on the intersections of a set of rows and columns each time. While quadratically reducing controls, it may necessitate more depth. We formulate the depth-optimal rectangular addressing problem as exact binary matrix factorization, an NP-hard problem also appearing in communication complexity and combinatorial optimization. We introduce a satisfiability modulo theories-based solver for this problem, and a heuristic, row packing, performing close to the optimal solver on various benchmarks. Furthermore, we discuss rectangular addressing in the context of fault-tolerant quantum computing, leveraging a natural two-level structure.
@@ I Introduction
Our motivation stems from recent successful large-scale experiments on the neutral atom arrays platform [1] that highlights the effect of reducing control complexity. As depicted in Figure 1a, the acousto-optic deflector (AOD) illuminates a product of rows and columns. Quantum gates, induced by specific pulses modulated by the AOD, address qubits at the row and column intersections. AODs prove effective for implementing gates [2, 1] and qubit movements [3, 4, 5, 6].
For a 2D array $X\times Y$, a (combinatorial) rectangle is a set of the form $X^{\prime}\times Y^{\prime}$, where $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$. Specifying a rectangle requires $|X|+|Y|$ bits (one bit for each row and each column), a significant reduction compared to $|X|\cdot|Y|$ bits for all elements. This quadratic reduction is maintained while preserving individual addressability, as a single element can still be considered a rectangle.
The coarser granularity of rectangular addressing may reduce control complexity at the cost of increasing depth. A generic example problem is given by the matrix in Figure 1b, where the qubits to address are represented by the 1’s. This matrix can be partitioned into five rectangles, each designated by distinct markers. Consecutively, each rectangle receives a modulated $R_{z}$ pulse through specific AOD configurations. Minimizing the number of rectangles to partition arbitrary binary matrices becomes crucial.
@@
In this paper, we consider the problem of achieving depth-optimal rectangular addressing, which we formulate as exact binary matrix factorization. We present a satisfiability modulo theories (SMT, extension of SAT) formulation for this problem and an effective heuristic dubbed row packing. The combined algorithm, SAP (SMT and packing), finds high-quality heuristic solutions quickly and then iteratively approaches to the optimal solution.
To assess our methods, we generate three benchmark sets. The first set comprises random matrices, the second includes matrices with known optimal solutions, despite the problem generally being NP-hard. The third set is constructed to accentuate the gap between the partition number and the conventional matrix rank.
We also observe that for future fault-tolerant quantum computing (FTQC), the problem may exhibit a product structure. This implies that we can solve limited-size problems on multiple levels and then combine the solutions.